OpenStudy (anonymous):

unbiased coin is tossed four times. Find the probability of the given event. (Round your answer to four decimal places.) The coin lands heads more than once.

OpenStudy (anonymous):

there is a more or less easy way to do with with a formula, but your sample space has only \(2^4=16\) elements in it, and 16 is not such a big number you can list them all and count

OpenStudy (anonymous):

is it clear what i mean by "list them all"?

OpenStudy (anonymous):

yes it did. Thank you. The answer was 11/16

OpenStudy (anonymous):

Five balls are selected at random without replacement from an urn containing four white balls and six blue balls. Find the probability of the given event. (Round your answer to three decimal places.) All of the balls are blue.

OpenStudy (kropot72):

P(first ball blue) = 6/10 P(second ball blue) = 5/9 P(third ball blue) = 4/8 P(fourth ball blue) = 3/7 P(fifth ball blue) = 2/6 P(sixth ball blue) = 1/5 The probability that all the balls are blue is \[\frac{6}{10}\times\frac{5}{9}\times\frac{4}{8}\times\frac{3}{7}\times\frac{2}{6}\times\frac{1}{5}=you\ can\ calculate\]

OpenStudy (anonymous):

i calculated 1/210 but its saying its wrong

OpenStudy (kropot72):

Sorry, I didn't read the question properly. Only five balls were selected. The probability that all five are blue is \[P(5\ balls\ are\ blue)=\frac{6}{10}\times\frac{5}{9}\times\frac{4}{8}\times\frac{3}{7}\times\frac{2}{6}=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you. I got the right answer .024

OpenStudy (kropot72):

Correct. Good work :)

OpenStudy (anonymous):

A customer from Cavallaro's Fruit Stand picks a sample of 3 oranges at random from a crate containing 75 oranges, of which 4 are rotten. What is the probability that the sample contains 1 or more rotten oranges? (Round your answer to three decimal places.)